A rectangle has a perimeter of 100cm. Find the greatest possible area for the rectangle using quadratic functions.
Found 2 solutions by ikleyn, Clanther:
Answer by ikleyn(52803)

.

A rectangle with a given perimeter which has the maximal area is a square.'


See the lesson
    - A rectangle with a given perimeter which has the maximal area is a square
in this site.



Answer by Clanther(6)

2l+2w=100
2(25)+2w=100
50+2w=100
2w=100-50
2w/2=50/2
w=25
A=lw
=25(25)
A=625cm²